RV Lough Foyle LF 4002
Cruise summary report
|Ship name (ship code)||RV Lough Foyle (74LG)|
|Cruise identifier||LF 4002|
|Cruise period||2002-09-29 — 2002-10-01|
|Port of departure||Belfast, United Kingdom|
|Port of return||Belfast, United Kingdom|
1. To maintain a nutrient and remote monitoring programme at mooring stations38A and 47D.
2. To assess temperature, salinity and nutrient distributions over depth at stations 38A and 47D.
Sunday 29 September 2002. In preparation for the cruise, all DANI scientific crew were onboard by 1930 hrs when mooring components and the automated sampler were prepared for deployment. Following a talk on ship's safety and a demonstration of personal life saving equipment, the RV Lough Foyle departed Belfast at 2055 hrs and sailed overnight in a fresh south easterly wind towards mooring site station 38A.
Monday 30 September 2002. The vessel arrived on the mooring site at 0600 hrs. The weather was dry with a fresh south easterly breeze when work for the day started at 0745 hrs. Problems were immediately encountered during mooring recovery when the torroid buoy became inverted in the water. Skilful work by the crew of the Lough Foyle eventually righted the buoy and the complete instrument mooring was recovered to ship deck at 0830 hrs. The mooring components were inspected for corrosion and parts replaced where necessary. The thermistor chain was removed from the mooring wire, temperature data downloaded and individual units reprogrammed. The CTDs and fluorometer were also removed, data downloaded and reprogrammed. The sub surface water sampler was serviced; samples removed, rebuilt, reprogrammed and attached to the mooring wire. The mooring components were reassembled and readied for deployment. The CTDs and fluorometers were attached and the mooring deployed at 1305 hrs on position 530 46l .850N 050 38l .065W. Following the deployment of the rosette water sampler and zooplankton net the survey continued to the coastal mooring station 47D. On arrival the complete instrument mooring was recovered to ship deck at 1615 hrs. The mooring components were inspected for corrosion and parts replaced where necessary. The CTD and fluorometer were removed, data downloaded and reprogrammed. The mooring components, CTD and fluorometer were then reassembled and the mooring was successfully redeployed at 1819 hrs on position 530 44l .490N 060 04l .000W. The water and zooplankton sampling operations were then repeated. Work on the station was completed at 1930 hrs and the vessel sailed to dock in Belfast at 0800 hrs Tuesday morning.
Tuesday 1 October 2002. Work for the day commenced at 0800 hrs when equipment was dismantled and removed from the vessel for return transportation to Newforge Lane. The scientific crew disembarked at 1045 hrs.
|Chief scientist||Brian Stewart (Department of Agriculture and Rural Development for Northern Ireland)|
|Cruise report||(0.02 MB)|
|General||Irish Sea and St. George's Channel|
|Specific||Irish Sea (north); ICES div. VIIa|